In a box ,there are 3 red marbles,4green marbles and 5 yellow marbles.what is the probability of drawing three marbles if draw simultaneously
A.they are of the same color?
B.they ate of three different colors?

Answer by rothauserc(4718)

A. Probability(P) of 3 red is (3/12) * (2/11) * (1/10) = 0.0045
P of 3 green is (4/12) * (3/11) * (2/10) = 0.0182
P of 3 yellow is (5/12) * (4/11) * (3/10) = 0.0455
P of 3 red or 3 green or 3 yellow = 0.0045 + 0.0182 + 0.0455 = 0.0682
:
B. P of 3 different colors = (3/12) * (4/11) * (5/10) = 0.0455
: